Calculate the mass of hydrogen formed when 30 grams of aluminum reacts with excess hydrochloric acid according to the balanced equation below:

2Al + 6HCl --> 2AlCl3 + 3H2

1
Expert's answer
2021-05-03T07:27:09-0400

n = m / M

M (Al) = 27 g/mol

M (H2) = 2 g/mol

n (Al) = 30 / 27 = 1.1 mol

According to the equation, n (H2) = 3/2 x n (Al) = 3/2 x 1.1 = 1.7 mol

m (H2) = n x M = 1.7 x 2 = 3.4 g